# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

### v2

@@ IX Limitations
Although not the main focus of this work, the current form of low-level policy introduces several limitations. One limitation is the applicable type of terrains. Currently the low-level policy is only trained on flat terrain in simulation. Although it shows the potential to work across different types of grounds (e.g. concrete, tile floor, carpet, mulch etc.), it does not push the low-level policy learning to its full potential. We can potentially leverage the recent advances in locomotion policy training [62, 7] to further expand the ability and applicability of SLIM across more diverse types terrains, e.g. stairs, ramp etc.
Another low-level-related limitation is that the current form of low-level policy is not a wholebody policy. Although the current form of low-level policy is sufficient for the task considered in this work, by generalizing it to a wholebody policy [15], we can potentially further enlarge its potential workspaces.
As a pure sim-to-real approach that is trained from scratch, we did not leverage any pre-trained foundation models. As a result, we did not address large amount of visual concepts and language variations and instead focused on the right amount of variations that are both practical and representative enough.
There are also a few limitations on the robot system side. Currently for the visual perception, we only used the RGB image from the wrist-mounted camera, which is the minimum setting that is sufficient for the current work. To further improve the robot’s perception and therefore the overall ability in terms of navigation and obstacle avoidance, it might be necessary to incorporate additional sensors for perception such as depth, stereo and lidar etc. In additiona, this can also provide valuable information for the robot to improve its longer-term memories via SLAM-like mechanisms [57, 5].
@@ X Conclusion, Discussion and Future Work
We present a complete legged mobile manipulation robot system for solving language-instructed long horizon task in this work. The policy is trained via the sim-to-real paradigm, together with a progressive policy expansion-based teacher policy for solving the long-horizon task, followed with a distillation-guided RL approach for student visuomotor policy learning. We further identified and reduced a number of crucial sim-to-real gaps. Experimental results compared with a number of baseline methods verified the effectiveness of the proposed SLIM method. Moreover, real-world testing of SLIM across different scenes and spatial layouts shows that the proposed method can generalize well, further demonstrating the effectiveness of the proposed whole pipeline.
@@ A-A Subtask Boundaries, Rewards and Behavior Priors
Since we decompose the full long task into a sequence of subtasks, in practice, we need to decide the task boundaries in order to properly divide the full task. In the sequential decompositional case as described in Eqn.(6), the task boundaries $\tau^{i}|\tau^{i+1}$ between two adjacent subtasks $\tau^{i}\rightarrow\tau^{i+1}$ can be defined by the success condition $\mathcal{S}(\tau^{i})$ of the subtask $\tau^{i}$ (i.e. $\tau^{i}|\tau^{i+1}\equiv\mathcal{S}(\tau^{i})$), since the success of a subtask $\tau^{i}$ naturally leads to the subsequent subtask $\tau^{i+1}$.
@@
    $\mathcal{S}$(SearchWObj): succeeds when the target object is in gripper and the target basket enters the view of the wrist-mounted RGB camera;
@@
    $\mathcal{S}$(MoveToWObj): succeeds when the target object is in gripper and the target basket enters a workspace of the robot within arm’s reach;
@@
    $\mathcal{S}$(MoveGripperToWObj): succeeds when the target object is in gripper and the grippers moved to the top of the target basket;
@@
We summarize the description of the subtask reward in Table II. The task reward represents the sparse task success ($+1$) or failure ($-1$) reward. For timeout, we used a reward of 00. Apart from the sparse task reward, we also incorporated some shaping rewards as shown in Table II. Most of the shaping rewards are delta-distance based shaping reward, encouraging two position of interest to getting closer to each other. Arm retract reward is also an instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (c.f. Figure A.13). Intuitively, the arm retract reward will encourage the robot’s arm to raise up after grasping the object and move close to its neutral position. Alignment reward is also an instance is delta-distance based shaping reward, where the distance is defined based on the cosine between the left and right finger-tip-to-object vectors. Keep-grasping reward is a contact based reward issuing a penalty of $-0.1$ is the contact between gripper fingers and the object is lost.
TABLE II: Rewards and Behavior Prior for Subtasks
@@
SearchWObj task + arm-retract rotational
MoveToWObj task + distance + arm-retract + keep-grasping
MoveGripperToWObj task + distance + keep-grasping stationary
@@
  • rotational: the forward velocity command ($\mathbf{c}[0]=0$) to the quadruped is kept as zero and therefore enable the quadruped to move in a rotational fashion. This is used as the behavior prior for search related subtasks (Search, SearchWObj).
  • stationary: both the forward and angular velocity of the command to the quadruped are hardcoded as zero ($\mathbf{c}=0$). This encourages the robot to be stationary. It is used in manipulation related tasks (Grasp, MoveGripperToWObj, DropInto).
In practice, both priors are implemented in a similar way, by initializing the teacher’s actor network for those subtasks with these priors to be of zero mean and small std for the corresponding action dimensions (e.g. the forward velocity dimension for rotational prior and both the forward and angular velocity dimensions for stationary prior). The reason for this design is to accommodate the KL-based distillation step in student learning, by providing a valid action distribution and being consistent with subtasks without these priors.
Hierarchical RL.
@@
Following this structure and given a frozen low-level policy, we are interested in training the high-level (teacher/student) policies. Because of this, the RL is essentially running at the high-level. Therefore, the per-(low-level-)step reward during each re-planning period is accumulated as the step reward for one high-level step, i.e.
@@
A-B Low-level Task and Reward
@@
Details of low-level training tasks/commands and corresponding rewards are summarized in Table III.
TABLE III: Low-level Training Tasks and Rewards
@@
gait following (optional)
@@
The domain randomization setting for low-level training is summarized in Table IV.
TABLE IV: Low-level Domain Randomization Parameters
@@ Appendix B Model Details
B-A Teacher High Level
The teacher network $\Pi\!=\!\{\pi^{k}\}_{k=1}^{K}$ is a set of sub-policies (referred to as policy set), as shown in Figure 3. The effective size of the policy set is $K\!=\!8$ (including the Idle subtask).
@@ B-A Teacher High Level
B-B Student High Level
@@
An illustration of the U-Net [42] for predicting segmentation and depth maps is shown in FigureA.14. We will talk about the architecture details by input modality below.
@@
An illustration of the policy input encoder network is shown in Figure A.15. The network is responsible to digest predicted segmentation&depth maps and proprioceptive observations, and produce a compact representational embedding for downstream RL networks. Again, we use a temporal stack of $N=5$ for both segmentation and depth map inputs. After converting each segmentation map to a one-hot representation, we stack all maps along the channel dimension, resulting in an input of $5\times(4+1)=25$ channels. The input is fed to a ResNet [21] of 5 residual blocks, each block with 64 filters, a kernel size of $3\times 3$, and a stride of 2. The output from the ResNet is spatially flattened, resulting in a sequence of vision tokens. On the other hand, we use an MLP with one hidden layer of 256 units to project the concatenated proprioceptive observation into a token embedding of 256 dimensions. This proprioceptive token is appended to the vision token sequence which is fed to a representation transformer with self-attention. The transformer has 5 layers, each layer with 8 attention heads. Each attention head has a key dimensionality of 128 and a value dimensionality of 128. The MLP for generating the residual output has 256 hidden units. Finally, we take the transformed output corresponding to a special learnable CLS token as the final output of the entire policy input encoder.
@@
The teacher policy incorporates a stationary behavior prior into the locomotion command $\mathbf{c}$ for several subtasks (Table II). With this prior, the corresponding locomotion command distribution always has a zero mean and a small std. Taking argmax of the distribution, the teacher policy generates exactly zero locomotion command and thus the quadruped can remain stationary. However, if we directly distill this distribution into the student’s counterpart, it is almost impossible to guarantee a zero mode, due to the continuous nature of the distribution. Thus we employed a sationary bit trick to solve this issue. Specifically, we create an extra binary action distribution for the student for predicting whether the quadruped should be stationary or not, which is determined by checking if the locomotion command of the teacher is exactly zero. We then add a new prediction loss for learning this stationary bit, while masking out the KLD loss (only for the two locomotion command dimensions) if the stationary bit is true. For student inference, whenever the stationary bit distribution outputs 1, the locomotion command $\mathbf{c}$ will be overwritten to be zero.
@@ B-C Low Level
For the low-level training, we used PPO [44] with a regularized online adaptation module Fu et al. [15].
@@
As shown in Figure 2, the low-level locomotion policy takes two types of inputs: 1) the standard proprioceptive state $\mathbf{s}_{\textrm{leg}}$ obtained from the environment, 2) the locomotion command $\mathbf{c}$ from an external command issuer (e.g.. The second type of input is crucial to make the low-level policy useful in downstream tasks where locomotion is involved as a basic skill for exploration and learning. During training, the command is re-sampled periodically within each episode and the low-level policy is trained to follow the command $\mathbf{c}$ via the command-following rewards detailed in Table III. The arm pose is also adjusted periodically within each episode, to make the low-level policy’s command following robust against arm movement. For the standard proprioceptive input to the low-level policy, while it is also possible to include those from the arm. Empirically, we discover that it might trigger some shakiness behavior in certain cases and therefore we only used $\mathbf{s}_{\textrm{leg}}$.
After training, the low-level policy can be used to follow commands issued by an external command issuer, e.g. either a higher level policy (as in SLIM and all the other baseline except Human Teleop), or human tele-operator (as in Human Teleop) in this work.
@@
. Similar as in [15, Figure 2], our online adaptation module consists of two encoders, the privileged encoder and the adaptation encoder. The privileged encoder takes privileged information as input and predicts a latent representation of environment extrinsics, while the adaptation encoder predicts the same latent representation depending only on robot sensory observations. Both encoders are trained to match each other and jointly with the RL objective. The privileged encoder network has two hidden layers of $(128,64)$ and the adaptation encoder has hidden layers of $(256,128)$.
B-D Human Teleoperation
To show that we’ve made reasonable efforts to make Human Teleop work, we list its failure modes. 1) The most common failure mode is task timeout. This is sometimes due to not fully grasping the cube and the cube slipping out of the fingers during movement, leaving no time to search and re-grasp. Failing to grasp fully can be due to either stopping a bit far from the cube, or incorrectly assessing the distance to the cube under monocular vision. 2) The arm can enter a twisted pose due to poor IK solutions, despite having already tuned IK as we can.
Human teleoperation is in general less accurate and slower than RL policy, due to Low level policy shaking, and human operator fatigue. Besides being slower, human teleoperation seems to be less energy efficient than our RL policies. When RL learned high-level policies can run smoothly on a particular low level policy, human teleoperation on the same low level policy seems to trigger over-current protection of Go1. This often happens when trying to raise an over-extended arm, causing Go1 to lose control of its leg motors. In evaluations, we ignore these Go1 hardware failures, and let the human operator retry for free.
B-E Model Parameter Count
We summarize in Table V the rough number of model parameters for different components of our system, including low level, teacher high level and student high level.
TABLE V: Number of Trainable Parameters.
@@ Appendix C Important Training Hyperparameters
Appendix D More Sim2Real Techniques
D-A More Dynamics Gap Reduction Techniques
Minimizing Tracking Error via PID Control.

### v3

@@ IX Limitations
Although not the main focus of this work, the current form of low-level policy introduces several limitations. One limitation is the applicable type of terrains. Currently the low-level policy is only trained on flat terrain in simulation. Although it shows the potential to work across different types of grounds (e.g., concrete, tile floor, carpet, mulch, etc.), it does not push the low-level policy learning to its full potential. We can potentially leverage the recent advances in locomotion policy training [67, 8] to further expand the ability and applicability of SLIM across more diverse types terrains, e.g., stairs, ramp, etc.
Another low-level related limitation is that the current form of low-level policy is not a wholebody policy. Although the current form of low-level policy is sufficient for the task considered in this work, by generalizing it to a wholebody policy [16], we can further enlarge its potential workspace.
As a pure sim-to-real approach that is trained from scratch, we did not leverage any pre-trained foundation models. As a result, we did not address a large amount of visual concepts and language variations and instead focused on the right amount of variations that are both practical and representative enough.
There are also a few limitations on the robot system side. Currently for the visual perception, we only used the RGB image from the wrist-mounted camera, which is the minimum setting that is sufficient for the current work. To further improve the robot’s perception and therefore the overall ability in terms of navigation and obstacle avoidance, it might be necessary to incorporate additional sensors for perception such as depth, stereo, lidar etc. In addition, this can also provide valuable information for the robot to improve its longer-term memories via SLAM-like mechanisms [63, 7].
@@ X Conclusion, Discussion and Future Work
We present a complete legged mobile manipulation robot system for solving language-instructed long-horizon tasks in this work. The policy is trained via the sim-to-real paradigm, together with a progressive policy expansion-based teacher policy for solving the long-horizon task, followed with a distillation-guided RL approach for student visuomotor policy learning. We further identified and designed a suite of crucial techniques for reducing sim-to-real gap. Experimental results compared with a number of baseline methods verified the effectiveness of the proposed SLIM method. Moreover, real-world testing of SLIM across different scenes and spatial layouts shows that the proposed method can generalize well, further demonstrating the effectiveness of the whole system.
@@ A-A Subtask Boundaries, Rewards and Behavior Priors
Since we decompose the full long task into a sequence of subtasks, in practice, we need to decide the task boundaries in order to properly divide the full task. In the sequential decompositional case as described in Section VI, the task boundaries $\tau^{i}|\tau^{i+1}$ between two adjacent subtasks ($\tau^{i}\rightarrow\tau^{i+1}$) can be defined by the success condition $\mathcal{S}(\tau^{i})$ of the subtask $\tau^{i}$ (i.e., $\tau^{i}|\tau^{i+1}\equiv\mathcal{S}(\tau^{i})$), since the success of a subtask $\tau^{i}$ naturally leads to the subsequent subtask $\tau^{i+1}$.
@@
    $\mathcal{S}$(SearchWithObj): succeeds when the target object is in gripper and the target basket enters the view of the wrist-mounted RGB camera;
@@
    $\mathcal{S}$(MoveToWithObj): succeeds when the target object is in gripper and the target basket enters a workspace of the robot within arm’s reach;
@@
    $\mathcal{S}$(MoveGripperToWithObj): succeeds when the target object is in gripper and the grippers moved to the top of the target basket;
@@
We summarize the description of the subtask reward in Table IV. The task reward represents the sparse task success ($+1$) or failure ($-1$) reward. For timeout, we used a reward of 00. Apart from the sparse task reward, we also incorporated some shaping rewards, as shown in Table IV. Most of the shaping rewards are delta-distance based shaping reward, encouraging two position of interest to getting closer to each other. Arm retract reward is also an instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (please refer to Fig. 6 for a reference of the neutral arm pose). Intuitively, the arm retract reward will encourage the robot’s arm to rise up after grasping the object and move close to its neutral position. Alignment reward is also an instance is delta-distance based shaping reward, where the distance is defined based on the cosine between the left and right finger-tip-to-object vectors. Keep-grasping reward is a contact based reward issuing a penalty of $-0.1$ is the contact between gripper fingers and the object is lost.
TABLE IV: Rewards and Behavior Prior for Subtasks
@@
SearchWithObj task + arm-retract rotational
MoveToWithObj task + distance + arm-retract + keep-grasping
MoveGripperToWithObj task + distance + keep-grasping stationary
@@
  • rotational: the forward velocity command ($\mathbf{c}[0]=0$) to the quadruped is kept as zero and therefore enable the quadruped to move in a rotational fashion. This is used as the behavior prior for search related subtasks (Search, SearchWithObj).
  • stationary: both the forward and angular velocity of the command to the quadruped are hardcoded as zero ($\mathbf{c}=0$). This encourages the robot to be stationary. It is used in manipulation related tasks (Grasp, MoveGripperToWithObj, DropInto).
In practice, both priors are implemented in a similar way, by initializing the teacher’s actor network for those subtasks with these priors to be of zero mean and small std for the corresponding action dimensions (e.g., the forward velocity dimension for rotational prior and both the forward and angular velocity dimensions for stationary prior). The reason for this design is to accommodate the KL-based distillation step in student learning, by providing a valid action distribution and being consistent with subtasks without these priors.
A-B Hierarchical RL and High-Level RL Training
@@
Following this structure and given a frozen low-level policy, we are interested in training the high-level (teacher/student) policies. Because of this, the RL is essentially running at the high-level. Therefore, the per-(low-level-)step reward during each re-planning period is accumulated as the step reward for one high-level step, i.e.,
@@
A-C Low-level Task and Reward
@@
Details of low-level training tasks, commands, and corresponding rewards are summarized in Table V.
Quadraped Gait Control.
Motivated by [39], we also introduce gait alignment rewards that incentivize the robot to follow predefined gait patterns. The gait patterns are defined by a phase offset vector of length three, where each entry represents the phase difference of each of the three legs (excluding the first one) relative to the first leg (fore left). The range of each entry is $[-0.5,0.5]$, where 00 indicates a synchronous movement with the first leg and $-0.5$ or $0.5$ indicates a half cycle phase difference. The phase of a leg is evaluated independently for each leg, based on the leg’s schedule at current time. The schedule is a number in $[0,1]$. For gait alignment rewards design, schedule $1$ represents the “swing” phase where smaller movements in the X-Y plane will be penalized; schedule 00 represents the “contact” phase where lower contact force will be penalized; for any schedule between $(0,1)$, both aspects will be penalized with weights modulated by the schedule value.
For training of the gait following ability, we expand the input command $\mathbf{c}$ with randomly sampled phase offset vector $\mathbf{g}_{\rm offset}$ and gait cycle vector $\mathbf{g}_{\rm cycle}$. $\mathbf{g}_{\rm offset}$ is sampled from five predefined gait patterns, as summarized in Table VI. $\mathbf{g}_{\rm cycle}$ is a two-dimensional vector, with each dimension representing the frequency and the stance ratio respectively. The frequency takes values in $[2.0,4.0]$, indicating the number of cycles that each leg goes through per second. The stance ratio takes values in $(0.4,0.6)$, indicating the portion of time within a cycle when each foot is in contact with the floor.
After low-level training, we fix the gait pattern to be “trot”, i.e., $\mathbf{g}_{\rm offset}\!=\![-0.5,-0.5,0.0]$, and the gait cycle vector as $\mathbf{g}_{\rm cycle}\!=\![2.4,0.5]$. We have found that this gait is suitable for the mobile manipulation task tackled in this work in terms of movement efficiency and stability. For the stationary task of low-level training, we always set $\mathbf{g}_{\rm cycle}\!=\![0,1]$.
TABLE V: Low-level Training Tasks and Rewards
@@
TABLE VI: Predefined Gait Patterns
Gait Names Phase Offsets
pronk $(0.0,0.0,0.0)$
walk $(-0.5,0.25,-0.25)$
trot $(-0.5,-0.5,0.0)$
bound $(0.0,0.5,0.5)$
pace $(-0.5,0.0,-0.5)$
@@
The domain randomization setting for low-level training is summarized in Table VII.
TABLE VII: Low-level Domain Randomization Parameters
@@ Appendix B Model Details
B-A Teacher (High Level)
The teacher network $\Pi\!=\!\{\pi^{k}\}_{k=1}^{K}$ is a set of sub-policies (referred to as policy set), as shown in Fig. 3. The effective size of the policy set is $K\!=\!8$ (including the Idle subtask).
@@ B-A Teacher (High Level)
Figure A.11: The architecture of the student’s policy input encoder that generates a compact observation embedding for RL.
B-B Student (High Level)
@@
An illustration of the U-Net [45] for predicting segmentation and depth maps is shown in Fig.A.10. The architecture details will presented according to input modality below.
@@
An illustration of the policy input encoder network is shown in Fig. A.11. The network is responsible to digest predicted segmentation&depth maps and proprioceptive observations, and produce a compact representational embedding for downstream RL networks. Again, we use a temporal stack of $N=5$ for both segmentation and depth map inputs. After converting each segmentation map to a one-hot representation, we stack all maps along the channel dimension, resulting in an input of $5\times(4+1)=25$ channels. The input is fed to a ResNet [23] of 5 residual blocks, each block with 64 filters, a kernel size of $3\times 3$, and a stride of 2. The output from the ResNet is spatially flattened, resulting in a sequence of vision tokens. On the other hand, we use an MLP with one hidden layer of 256 units to project the concatenated proprioceptive observation into a token embedding of 256 dimensions. This proprioceptive token is appended to the vision token sequence which is fed to a representation transformer with self-attention. The transformer has 5 layers, each layer with 8 attention heads. Each attention head has a key dimensionality of 128 and a value dimensionality of 128. The MLP for generating the residual output has 256 hidden units. Finally, we take the transformed output corresponding to a special learnable CLS token as the final output of the entire policy input encoder.
@@
The teacher policy incorporates a stationary behavior prior into the locomotion command $\mathbf{c}$ for several subtasks (Table IV). With this prior, the corresponding locomotion command distribution always has a zero mean and a small std. Taking argmax of the distribution, the teacher policy generates exactly zero locomotion command and thus the quadruped can remain stationary. However, if we directly distill this distribution into the student’s counterpart, it is almost impossible to guarantee a zero mode, due to the continuous nature of the distribution. Thus we employed a sationary bit trick to solve this issue. Specifically, we create an extra binary action distribution for the student for predicting whether the quadruped should be stationary or not, which is determined by checking if the locomotion command of the teacher is exactly zero. We then add a new prediction loss for learning this stationary bit, while masking out the KLD loss (only for the two locomotion command dimensions) if the stationary bit is true. For student inference, whenever the stationary bit distribution outputs 1, the locomotion command $\mathbf{c}$ will be overwritten to be zero.
@@ B-C Low Level
For low-level training, we use PPO [47] with a regularized online adaptation module [16].
@@
As shown in Fig. 2, the low-level locomotion policy takes two types of inputs: 1) the standard proprioceptive state $\mathbf{s}_{\textrm{leg}}$ obtained from the environment, 2) the locomotion command $\mathbf{c}$ from an external command issuer (e.g., The second type of input is crucial to make the low-level policy useful in downstream tasks where locomotion is involved as a basic skill for exploration and learning. During training, the command is re-sampled periodically within each episode and the low-level policy is trained to follow the command $\mathbf{c}$ via the command-following rewards detailed in Table V. The arm pose is also adjusted periodically within each episode, to make sure the low-level policy can follow locomotion commands while being robust against arm movements. For the standard proprioceptive input to the low-level policy, while it is also possible to include those from the arm, empirically, we discover that it might cause the quadruped to shake more often, and decide to only use $\mathbf{s}_{\textrm{leg}}$. We also discover that the joint velocities in the hardware are delayed more than joint position readings from the motor sensors, possibly due to temporal smoothing. Since we do not know the exact algorithm that computes the joint velocities, we simply do not use them as input, and rely on feeding the neural networks the previous few frames for the policy to estimate the joint velocities.
After training, the low-level policy can be used to follow commands issued by an external command issuer. This can be either a higher level policy (as in SLIM or all the autonomous baselines), or human teleoperation (Human Teleop) in this work.
@@
. Similar as in [16, Fig.2], our online adaptation module consists of two encoders, the privileged encoder and the adaptation encoder. The privileged encoder takes privileged information as input and predicts a latent representation of environment extrinsics, while the adaptation encoder predicts the same latent representation depending only on robot sensory observations. Both encoders are trained to match each other and jointly with the RL objective. The privileged encoder network has two hidden layers of $(128,64)$ and the adaptation encoder has hidden layers of $(256,128)$.
B-D Model Parameter Count
We summarize in Table VIII the rough number of model parameters for different components of our system, including low level, teacher high level and student high level.
TABLE VIII: Number of Trainable Parameters.
@@ Appendix C Important Training Hyperparameters
Appendix D Simulator Choice
All simulations were done in the MuJoCo simulator [51]. At the beginning of the project, we evaluated several popular simulators at the time, including MuJoCo [^4] https://mujoco.org/, Isaac Gym [^5] https://developer.nvidia.com/isaac-gym, and PyBullet [^6] https://pybullet.org/wordpress/. We settled on MuJoCo due to its reasonably fast simulation (on CPU) and rendering (on GPU) at the same time.
Besides camera observation rendering for student training, due to partial observability, we also rely on segmentation rendering to check which objects are in the camera view and use it as privileged information for the teacher, during both teacher and student training. Thus RGB and segmentation rendering efficiencies are critical to our simulator choice.
Appendix E Sim-to-Real Techniques
E-A Dynamics Gap Reduction Techniques
The dynamics sim-to-real gap is the mismatch between the transition function $P(\mathbf{s}^{t}|\mathbf{s}^{t-1},\mathbf{a}^{t-1})$ of the simulation versus that of the real world. It can be due to the mismatch in the physical properties of objects and motors, or by things that are not properly simulated, such as friction and backlash. This difference in dynamics can cause policies trained purely in simulation to fail when deployed in the real world, especially if the task requires accurate motor control, such as grasping a small object. We use the following key techniques to address the dynamics gap.
Arm PID Control to Minimize Tracking Errors.